How much water must be added to a 20 quarts of a 12% alcohol solution to dilute it to a 9% solution?
Found 2 solutions by stanbon, modi.kk91:
Answer by stanbon(75887)

How much water must be added to a 20 quarts of a 12% alcohol solution to dilute it to a 9% solution?
-----
Equation:
water + water = water
0.88*20 + 1.00x = 0.91(20+x)
----
88*20 + 100x = 91*20 + 91x
-------
9x = 3*20
x = (1/3)20
x = 6 2/3 quarts (amt. of water to add)
=============
Cheers,
Stan H.

Answer by modi.kk91(20)

by allegation and mixture formula

o% 12%
9%
12-9=3 and 9-0=9
ratio=3:9=1:3
1:3::x:20

x=20/3